OpenStudy (jokersmart):

Check my answer? Please help out! :( What is the exact volume of the sphere? A. 56.3 ft3 B. 274.625 ft3 C. 366.16 ft3 D. 1464.6 ft3 <---

OpenStudy (igreen):

Picture? We need more info. @JokerSmart

OpenStudy (jokersmart):

Ah, yes sorry about that!

OpenStudy (igreen):

No, that's incorrect.

OpenStudy (igreen):

It wants the exact Volume, which is in terms of pi.

OpenStudy (igreen):

\(\sf V = \dfrac{4}{3} \pi r^3\) Take pi out: \(\sf V = \dfrac{4}{3}r^3\) \(\sf V = \dfrac{4}{3}(6.5)^3\) Simplify

OpenStudy (jokersmart):

Oh. Would it be C then? (second answer.) I kinda doubt it would be A...

OpenStudy (igreen):

Yep, that's what I got.

OpenStudy (igreen):

All of them have pi next to it, right?

OpenStudy (jokersmart):

Yeah ^-^

OpenStudy (jokersmart):

Thank you for your help! I appreciate it.